# Enabling Intelligent Interactions between an Agent and an LLM: A Reinforcement Learning Approach

### v6

@@ Abstract
Large language models (LLMs) encode a vast amount of world knowledge acquired from massive text datasets. Recent studies have demonstrated that LLMs can assist an embodied agent in solving complex sequential decision making tasks by providing high-level instructions. However, interactions with LLMs can be time-consuming. In many practical scenarios, they require a significant amount of storage space that can only be deployed on remote cloud server nodes. Additionally, using commercial LLMs can be costly since they may charge based on usage frequency. In this paper, we explore how to enable intelligent cost-effective interactions between the agent and an LLM. We find that this problem can be naturally formulated by a Markov decision process (MDP), and propose When2Ask, a reinforcement learning based approach that learns when it is necessary to query LLMs for high-level instructions to accomplish a target task. Experiments on MiniGrid and Habitat environments that entail planning sub-goals demonstrate that When2Ask learns to solve target tasks with only a few necessary interactions with an LLM, and significantly reduces interaction costs in testing environments compared with baseline methods. Experiment results also suggest that by learning a mediator model to interact with the LLM, the agent’s performance becomes more robust against partial observability of the environment. Our code is available at https://github.com/ZJLAB-AMMI/LLM4RL.
@@ 1 Introduction
To empower embodied agents with the capability to effectively handle demanding sequential decision-making tasks, it is essential for them to possess reasoning abilities that enable them to plan for the long-term consequences of their actions Deitke et al. (2022). Reinforcement learning (RL), particularly deep RL, has emerged as a popular paradigm for addressing these challenges. Deep RL involves agents interacting with the environment and learning from feedback to improve their decision-making over time. Despite recent advancements in deep RL, several challenges still remains and limits its vast applications in real world scenarios. For instance, solving complex problems using deep RL often requires significant computational resources. Additionally, safety concerns can arise during the learning phase, especially in scenarios where the agent’s exploration might interact with the real world or other sensitive environments Das et al. (2018); Chevalier-Boisvert et al. (2018). As an alternative, the emergence of large language models (LLMs) has shown promise in tackling these issues. Previous studies have demonstrated that LLMs possess reasoning capabilities Radford et al. (2019); Brown et al. (2020); Wei et al. (2022). Researchers have explored leveraging LLMs’ reasoning abilities to solve various embodied tasks, including robot manipulation tasks Ahn et al. (2022); Huang et al. (2022); Jiang et al. (2022) and playing video games Dasgupta et al. (2023); Wang et al. (2023a; c). As depicted in Fig. 1, the embodied agent interacts with the environment, gathering information ralated to the target task, and utilizes LLMs as explicit reasoners to make high-level plans using natural language instructions, such as instructing a robot to “pick up a can of coke” or “place an apple on the table” for the next step.
@@
While the integration of pre-trained LLMs as explicit planners in embodied agents has demonstrated promising results, enabling efficient interaction between these agents and LLMs to solve real-world problems remains challenging. Frequent queries to LLMs can result in unnecessary resource wastage, including fees (if a commercial LLM is used), communication overhead and reasoning time. Whereas insufficient queries to LLMs prevent the agent from adjusting its plan according to online feedback from the environment, making it vulnerable to uncertainties in the environment.
@@
Determining an appropriate guideline for querying LLMs requires expert knowledge of the target task. Consider a scenario where a robot is instructed to collect a can of coke but encounters a locked door on its way to the kitchen. Ideally, the agent should recognize this incident and adjust its plan accordingly by consulting the LLM on how to deal with the locked door. In such cases, timely decision-making regarding when to consult the LLM planner becomes crucial. Failure to interrupt the ongoing action plan and request a new one in time can hinder task completion progress or even lead to safety issues, such as damaging the door or the robot itself. Conversely, frequent requests for plans from the LLM can be time-consuming and costly, particularly when using commercial LLMs deployed on remote cloud server nodes that charge based on usage frequency.
@@
In this paper, we propose When2Ask, a general approach that trains the agent to make intelligent cost-effective interactions between itself and an LLM deployed on a remote cloud server. Our objective is to facilitate effective completion of a target task while minimizing communication costs incurred from interactions with the LLM. Specifically, we adopt a Planner-Actor-Mediator framework, similar to Dasgupta et al. (2023), where the planner is a pre-trained LLM used for making plans, the actor contains policies for executing the plans, and the mediator serves as an interface in between by deciding when to request a new plan and generating observation representations for the planner (which are text descriptions). With a focus on optimizing interacting timings, we use RL to learn an asking policy that instructs the agent to either adhere to the current plan or request a new plan from the LLM.
To summarize, our main contributions are as follows:
@@
  • We propose an RL approach termed When2Ask to coordinate the interaction between the agent and the LLM based on the Planner-Actor-Mediator framework Dasgupta et al. (2023). Concretely, we propose to introduce an explicit asking policy in the mediator and train it using an RL approach to determine when to query the LLM planner.
@@
  • We conducted a comprehensive evaluation of When2Ask against baseline methods based on simulation platforms MiniGrid Chevalier-Boisvert et al. (2023) and Habitat Szot et al. (2021). The results demonstrate that the learned asking policy is able to make intelligent decisions on when to query LLMs, resulting in high success rates with only a few necessary LLM interactions in test tasks. Additionally, we find that our approach can perform robustly against partial observability of the environments.
@@ 2.1 The Options Framework
We consider sequential decision-making in embodied environments, which is commonly formalized as a Markov decision process (MDP), denoted as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},p,r,\gamma\rangle$. Here $\mathcal{S}$ represents the state space, $\mathcal{A}$ represents the action spaces, $p(s^{\prime}|s,a)$ denotes the state transition probability function, $r(s,a)$ represents the reward function, and $\gamma$ is the discount factor. The objective in such a framework is to learn an optimal policy that maximizes the cumulative return over time $\sum_{t}\gamma^{t}r(s_{t},a_{t})$, where $t$ denotes the time index.
The options framework extends the traditional notion of action in an MDP to include options, which are essentially closed-loop policies that encompass a sequence of actions over a period of time Sutton et al. (1999); Precup (2000). Options can range from higher-level tasks such as picking up an object or going to lunch, to more primitive actions like muscle twitches and joint torques. The introduction of options allows for the incorporation of temporally abstract knowledge and action within the RL framework in a natural and general manner, thus provides a flexible and intuitive approach to handle complex tasks with varying levels of granularity.
Formally, an option $\omega$ is defined as a 3-tuple $\langle\mathcal{I_{\omega}},\pi_{\omega},\beta_{\omega}\rangle,$ where $\mathcal{I}_{\omega}$ represents the initial state set for this option, $\pi_{\omega}$ denotes the acting policy for the option, and $\beta_{\omega}$ represents the termination condition. Given a state $s$, a policy-over-options would select an option $\omega$ from the set of available options $\Omega$. The agent would then plan low-level actions by following its current option policy $a\sim\pi(\cdot|s,\omega)$ until the termination condition $\beta_{\omega}$ is satisfied. In our work, we use pre-defined skills as options and a pre-trained LLM as the policy-over-options to generate high-level options.
@@ 3 Related work
LLMs have emerged as powerful tools for plan generation, and previous research has focused on designing effective interfaces between planners and actors. In Ahn et al. (2022), LLMs are employed to plan the entire sequence of options at the beginning of each task, enabling the agent to complete the task without further interaction with the planner. In Wang et al. (2023c), the authors introduce a feedback system where the agent requests the LLM to generate an updated plan based on environmental feedback when the execution of the previous plan fails. This approach enhances the robustness of the acting agent in the face of environmental uncertainties. However, these methods often rely on hard-coded failure detectors, such as applying a threshold to limit the number of permissible MDP state-transition timesteps for an option. In Ren et al. (2023), a framework is proposed for measuring and aligning the uncertainty of LLM-based planners, allowing them to seek assistance from humans when necessary. In addition, Dasgupta et al. (2023) introduce the Planner-Actor-Reporter framework, which includes a reporter module to enhance information exchange between the actor and the LLM-based planner. In this framework, the agent interacts with the LLM at each timestep, regardless of whether new information is acquired or not. While this approach eliminates the need for hard-coded termination conditions and reduces uncertainties during option execution, it leads to excessive resource consumption, especially when utilizing a large-scale and expensive LLM as the planner.
In this paper, we propose learning an interaction policy that enables the local agent to communicate with a remote LLM, and empirically demonstrate that our approach can overcome the limitations of previously mentioned hard-coded rule-based interaction protocols or protocols that entail querying the LLM at each time point.
@@ 4.1 The Planner-Actor-Mediator Framework
Mediator  In this work, our primary focus is on designing an intelligent mediator component within the Planner-Actor-Mediator framework. Our approach involves training an explicit asking policy using RL to determine when to interact with the planner. The mediator component consists of two sub-components: an asking policy that decides whether to request a new plan from the planner based on observations and the current option, and a translator module that converts observations into text descriptions readable by the LLM. Following Ahn et al. (2022); Carta et al. (2023), we assume the availability of an expert translator here. In our experiments, the translator is designed with two stages. Firstly, we extract the IDs of objects, such as keys, doors, and boxes, observed within the field of view of the local agent using the built-in interface of the simulation platform. Next, we input this information into our predefined template and output it to LLM in a fixed format. An example of the format can be seen in the green box of Fig.4. The translator can also be learned from data Wang et al. (2023c); Dasgupta et al. (2023).
@@ 4.2 Learning asking policy with RL
As mentioned earlier, interacting with the LLM can be costly. Ideally, the asking policy should be trained to enable the agent to request a new plan from the LLM only when it discovers new and informative observations. The expectation is that the LLM will provide a different plan in response to these new observations. To address this, we formulate the problem as an MDP, where the state includes information about the agent’s observation and current option in action. The action space consists of two actions: “Ask" and “Not Ask". In this formulation, the LLM planner is considered as part of the environment that can influence state transitions. The reward function encompasses both the task-related return, denoted as $r$, and an additional penalty term that penalizes unnecessary interactions. Specifically, when the asking policy decides to ask the LLM for a new plan, but the plan provided by the LLM remains the same as the current one, the agent incurs a penalty. This penalty encourages the asking policy to avoid unnecessary interactions and ensures that requesting a new plan is primarily motivated by the discovery of new informative observations.
@@
[…]
@@ 5 Experiments
Through experiments, we seek to address the following questions: can our agent effectively minimize interaction costs while maintaining a high target task completion rate, compared with baseline methods? Can our agent proactively seek assistance from an LLM in exploratory environments? Lastly, how does the performance of our agent compare to a baseline RL agent that does not utilize an LLM planner? The experimental results indicate that the answer to all of the aforementioned questions is yes. As a byproduct, we find that our approach is able to tolerate the imperfection of a crucial component, the translator, in the mediator module, which is used to convert observed images into textual descriptions (refer details to the Appendix). We employ two versions of the Vicuna model (Vicuna-7b and Vicuna-13b) Touvron et al. (2023) as LLM planners in our experiments.
@@ 5.1 Baselines
Hard-coded  The timing and conditions for requesting new instructions from LLMs are manually determined by human experts for each option Wang et al. (2023c). The agent will only request a new plan from the LLM planner when specific termination conditions for the option are met. These conditions involve a goal-finishing detector and a constraint on the maximum number of allowed timesteps. For example, let’s consider the option “go to the red door." The termination condition for this option specifies that the agent should reach the target door location or exceed 100 timesteps spent on this option. We argue that, adopting such hard-coded termination rules, the agent cannot fully utilize newly acquired information during option execution. Additionally, these hard-coded rules may be vulnerable to uncertainties embedded in other components of the framework.
@@
Never  The agent never interacts with the LLM. Instead, the policy-over-options (i.e., the planner) is learned using RL techniques based on data collected during interactions with the environment Sutton et al. (1999); Precup (2000). This means that the agent learns to make decisions and generate plans without actively querying the LLM in real-time decision-making. By comparing this method with other approaches, we can assess the contribution of using an LLM as the planner for embodied sequential decision-making tasks. This comparison helps evaluate the effectiveness and advantages of incorporating a pre-trained LLM into the planning process.
@@ 5.2 MiniGrid Experiments
The MiniGrid environment Chevalier-Boisvert et al. (2023) consists of a collection of 2D grid-world environments with goal-oriented tasks. In these environments, the agent must navigate within a 2D grid room and interact with specific objects to complete various tasks, such as “open the red door" or “put the green ball next to the yellow box".
One important characteristic of our MiniGrid environment is that the agent’s view range is limited. This means that the agent needs to explore the environment and gather useful information to plan its actions effectively. The environment returns observations in the form of a full grid, but with unexplored areas occluded, similar to the concept of “fog of war" in games like StarCraft. Technically, the observation returned by the environment has a shape of $o\in\mathbb{R}^{W\times H\times 4}$, where $W$ and $H$ represent the width and height of the grid, respectively. For an unexplored grid at location $[w,h]$, the observation returns the vector $[-1,-1,-1,-1]$. For an explored grid, the corresponding 4D vector contains information about the object ID, color ID, state ID (e.g., closed or locked for a door), and the agent’s direction ID (indicating the agent’s orientation if it is present at this location, or 4 otherwise). This design allows us to focus on the agent’s reasoning ability and exclude potential influences from factors like memorization. Fig. 3 provides an example of the environment setup in the SimpleDoorKey scenario.
@@
Figure 4: Example of the prefix prompt and one interaction for the ColoredDoorKey environment. Prefix prompt consists of task instruction and few-shot examples. In Chain-of-Thought-style prompts, we add inference processes within the examples. Note that these few-shot examples are only provided for grounding task knowledge and constraining the output formats of the LLM. We do not need to exhaustively enumerate all knowledge and rules to construct prompts, as a qualified LLM can do logical reasoning based on a limited number of prompts, then provide proper plans (instructions) that are adaptable to new scenarios encountered in the environment.
In our experiments, we focus on the task of opening a locked door in five distinct environments: SimpleDoorKey, KeyInBox, RandomBoxKey, ColoredDoorKey, and MovingObstacle. All of these environments are procedurally generated, i.e., the grid layout (including room size, key and door locations) is randomly determined each time the environment is reset. To evaluate generalization, a held-out test set consisting of 100 randomly selected seeds is predefined for each environment. Refer to the Appendix for more information about the environments.
We use the Vicuna-7b model for the SimpleDoorKey, KeyInBox, RandomBoxKey, and MovingObstacle environments, while for the more complex ColoredDoorKey environment we use the Vicuna-13b model. To enable interactions between the agent and the LLM planner. As demonstrated in previous work Min et al. (2022), language models like LLMs require carefully designed prompts and few-shot demonstrations to generalize to different tasks. In our experiments, we provide task instructions and few-shot examples as in-context prompts for each environment. These prompts serve to ground the task knowledge and guide the LLM’s understanding of the specific task requirements. Furthermore, for the challenging reasoning task in the ColoredDoorKey environment, we utilize Chain-of-Thought prompts proposed by Wei et al. (2022). These prompts help the LLM to deal with complex reasoning tasks specific to the ColoredDoorKey environment. The few-shot examples provided in the prompts are used to anchor the agent’s knowledge about the task, such as understanding that a door can only be unlocked with a key of the same color, and provide constraints on the output formats and guidelines for generating appropriate plans. Note that the LLM planner must reason about the target task using its embedded knowledge and generalization capabilities to adapt to different scenarios with varying objects and colors. Fig. 4 provides an example of the prefix prompts and an interaction example in the ColoredDoorKey environment. It shows how the LLM planner successfully generates a correct plan based on novel observations.

### v8

@@ Abstract
Large language models (LLMs) encode a vast amount of world knowledge acquired from massive text datasets. Recent studies have demonstrated that LLMs can assist an embodied agent in solving complex sequential decision making tasks by providing high-level instructions. However, interactions with LLMs can be time-consuming. In many practical scenarios, it requires a significant amount of storage space that can only be deployed on remote cloud servers. Additionally, using commercial LLMs can be costly since they may charge based on usage frequency. In this paper, we explore how to enable intelligent cost-effective interactions between a down stream task oriented agent and an LLM. We find that this problem can be naturally formulated by a Markov decision process (MDP), and propose When2Ask, a reinforcement learning based approach that learns when it is necessary to query LLMs for high-level instructions to accomplish a target task. On one side, When2Ask discourages unnecessary redundant interactions, while on the other side, it enables the agent to identify and follow useful instructions from the LLM. This enables the agent to halt an ongoing plan and transition to a more suitable one based on new environmental observations. Experiments on MiniGrid and Habitat environments that entail planning sub-goals demonstrate that When2Ask learns to solve target tasks with only a few necessary interactions with the LLM, significantly reducing interaction costs in testing environments compared with baseline methods. Our code is available at: https://github.com/ZJLAB-AMMI/LLM4RL.
@@ 1 Introduction
To empower embodied agents with the capability to effectively handle demanding sequential decision-making tasks, it is essential for them to possess reasoning abilities that enable them to plan for the long-term consequences of their actions Deitke et al. (2022). Reinforcement learning (RL), particularly deep RL, has emerged as a popular paradigm for addressing these challenges. Deep RL involves agents interacting with the environment and learning from feedback to improve their decision-making over time. Despite recent advancements, several challenges still remains and limits its vast applications in real world scenarios. For instance, solving complex problems using deep RL often requires significant computational resources. Additionally, safety concerns can arise during the learning phase, especially in scenarios where the agent’s exploration might interact with the real world or other sensitive environments Das et al. (2018); Chevalier-Boisvert et al. (2018). As an alternative, the emergence of large language models (LLMs) has shown promise in tackling these issues. Previous studies have demonstrated that LLMs possess reasoning capabilities Radford et al. (2019); Brown et al. (2020); Wei et al. (2022). Researchers have explored leveraging LLMs’ reasoning abilities to solve various embodied tasks, including robot manipulation tasks Ahn et al. (2022); Huang et al. (2022); Jiang et al. (2022) and playing video games Dasgupta et al. (2023); Wang et al. (2023a; c). As depicted in Fig. 1, the embodied agent interacts with the environment, gathering information ralated to the target task, and utilizes LLMs as explicit reasoners to make high-level plans using natural language instructions, such as instructing a robot to “pick up a can of coke” or “place an apple on the table”.
@@
While the integration of pre-trained LLMs as explicit planners in embodied agents has demonstrated promising results, enabling efficient interaction between these agents and LLMs to solve real-world problems remains challenging. Frequent queries to LLMs can result in unnecessary resource wastage, including fees (if a commercial LLM is used), communication overhead and reasoning time. Whereas insufficient queries to LLMs prevent the agent from obtaining useful instructions in time to adjust its plan to respond to the complex and changing environment.
@@
Determining an appropriate guideline for querying LLMs requires expert knowledge of the target task. Consider a scenario where a robot is instructed to collect a can of coke but encounters a locked door on its way to the kitchen. Ideally, the agent should recognize this incident and adjust its plan accordingly by consulting the LLM on how to deal with the locked door. In such cases, timely decision-making regarding when to consult the LLM planner becomes crucial. Failure to interrupt the ongoing action plan and request a new one in time can hinder task completion progress or even lead to safety issues, such as damaging the door or the robot itself. Conversely, frequent requests for plans from the LLM can be time-consuming and costly, particularly when using commercial LLMs deployed on remote cloud servers that charge based on usage frequency.
@@
In this paper, we propose When2Ask, a general approach that trains the agent to make intelligent cost-effective interactions between itself and an LLM remotely deployed. Our objective is to facilitate effective completion of a target task while reducing unnecessary non-informative interactions with the LLM. Specifically, we adopt a Planner-Actor-Mediator framework, similar to Dasgupta et al. (2023), where the planner is a pre-trained LLM used for making plans, the actor contains policies for executing the plans, and the mediator serves as an interface in between by deciding when to request a new plan and generating observation representations for the planner (which are text descriptions). With a focus on optimizing interacting timings, we use RL to learn an asking policy that instructs the agent to either adhere to the current plan or request a new plan from the LLM.
To summarize, our main contributions include:
@@
  • We propose an RL approach termed When2Ask to coordinate the interaction between a down stream task oriented agent and a pre-trained LLM based on the Planner-Actor-Mediator framework Dasgupta et al. (2023). Specifically, we propose to introduce an explicit asking policy in the mediator and train it using an RL approach to determine when to query the LLM planner.
@@
  • We conducted a comprehensive evaluation of When2Ask against baseline methods based on simulation platforms MiniGrid Chevalier-Boisvert et al. (2023) and Habitat Szot et al. (2021). The results demonstrate that the learned asking policy is able to make intelligent decisions on when to query LLMs, resulting in high success rates with only a few necessary LLM interactions in the testing phase. In contrast to conventional interaction strategies that rely on predetermined termination criteria, When2Ask offers a significant advantage by enabling the interruption of an ongoing plan in favor of a new plan that addresses emerging observations.
@@ 2.1 The Options Framework
We consider sequential decision-making in embodied environments, which is commonly formalized as a Markov decision process (MDP), denoted as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},p,r,\gamma\rangle$. Here $\mathcal{S}$ represents the state space, $\mathcal{A}$ represents the action spaces, $p(s^{\prime}|s,a)$ denotes the state transition probability function, $r(s,a)$ represents the reward function, and $\gamma$ is the discount factor. The objective is to learn an optimal policy that maximizes the cumulative reward over time $\sum_{t}\gamma^{t}r(s_{t},a_{t})$, where $t$ denotes the time index.
The options framework extends the traditional notion of action in an MDP to include options, which are essentially closed-loop policies that encompass a sequence of actions over a period of time Sutton et al. (1999); Precup (2000). Options can range from higher-level tasks such as picking up an object or going to lunch, to more primitive actions like muscle twitches and joint torques. The introduction of options allows for the incorporation of temporally abstract knowledge and action within the RL framework in a natural and general manner, thus provides a flexible and intuitive approach to handle complex tasks with varying levels of granularity. Formally, an option $\omega$ is defined as a 3-tuple $\langle\mathcal{I_{\omega}},\pi_{\omega},\beta_{\omega}\rangle,$ where $\mathcal{I}_{\omega}$ represents the initial state set, $\pi_{\omega}$ denotes the acting policy, and $\beta_{\omega}$ represents the termination condition for this option. Given a state $s$, a policy-over-options would select an option $\omega$ from the set of available options $\Omega$. The agent would then plan low-level actions by following its current option policy $a\sim\pi(\cdot|s,\omega)$ until the termination condition $\beta_{\omega}$ is satisfied. In our work, we use pre-defined skills as options and a pre-trained LLM as the policy-over-options to generate high-level options.
@@ 3 Related work
LLMs have emerged as powerful tools for plan generation. There are some research works focusing on designing effective interfaces between planners and actors. In Ahn et al. (2022), LLMs are employed to plan the entire sequence of options at the beginning of each task, enabling the agent to complete the task without further interaction with the planner. In Wang et al. (2023c), the authors introduce a feedback system where the agent requests the LLM to generate an updated plan based on environmental feedback when the execution of the previous plan fails. This approach enhances the robustness of the acting agent in the face of environmental uncertainties. However, these methods often rely on hard-coded failure detectors or apply a threshold to limit the number of permissible MDP state-transition timesteps for an option. In Ren et al. (2023), a framework is proposed for measuring and aligning the uncertainty of LLM-based planners, allowing them to seek assistance from humans when necessary. In addition, Dasgupta et al. (2023) introduce the Planner-Actor-Reporter framework, which includes a reporter module to enhance information exchange between the actor and the LLM-based planner. In this framework, the agent interacts with the LLM at each timestep, regardless of whether new information is acquired or not. While this approach eliminates the need for hard-coded termination conditions and reduces uncertainties during option execution, it leads to excessive resource consumption, especially when utilizing a large-scale and expensive LLM as the planner.
In this paper, we propose learning an interaction policy that enables the agent to interact with a remote LLM in an autonomous and “smarter” way. We empirically demonstrate that our approach overcomes the limitations of previously mentioned hard-coded rule-based interaction protocols or protocols that entail querying the LLM at each timestep.
@@ 4.1 The Planner-Actor-Mediator Framework
Mediator  In this work, our primary focus is on designing an intelligent mediator component within the Planner-Actor-Mediator framework. Our approach involves training an explicit asking policy using RL to determine when to interact with the planner. The mediator component consists of two sub-components: an asking policy that decides whether to request a new plan from the planner based on observations and the current option, and a translator module that converts observations into text descriptions readable by the LLM. Following Ahn et al. (2022); Carta et al. (2023), we assume the availability of an expert translator here. In our experiments, the translator is designed with two stages. Firstly, we extract the IDs of objects, such as keys, doors, and boxes, observed within the field of view of the agent using the built-in interface of the simulation platform. Next, we input this information into our predefined prompt template and output it to LLM in a fixed format. An example of the format can be seen in the green box of Fig.4. The translator can also be learned from data Wang et al. (2023c); Dasgupta et al. (2023).
@@ 4.2 Learning asking policy with RL
As mentioned earlier, interacting with the LLM can be costly. Ideally, the asking policy should be trained to enable the agent to request a new plan from the LLM only when it discovers new and informative observations. The expectation is that the LLM will provide a different plan in response to these new observations. To address this, we formulate the problem as an MDP, where the state includes information about the agent’s observation and current option in action. The action space consists of two actions: “Ask” and “Not Ask”. In this formulation, the LLM planner is considered as part of the environment that can influence state transitions. The reward function consists of both the task-related return, denoted as $r$, and an additional penalty term that penalizes unnecessary interactions. Specifically, when the asking policy decides to ask the LLM for a new plan, but the plan provided by the LLM remains the same as the current one, the agent incurs a penalty. This penalty encourages the asking policy to avoid unnecessary interactions and ensures that requesting a new plan is primarily motivated by the discovery of new informative observations.
@@
$\max_{\theta}\sum_{t=1}\big{[}\gamma^{t}r_{t}-\lambda\mathbbm{1}(y_{t}==\textit{Ask}\land\omega_{t}==\omega_{t-1})\big{]},$ (1)
where $y_{t}\in\{\textit{Ask},\textit{Not Ask}\}$ represents the decision made by the asking policy at timestep $t$, $r_{t}$ denotes the task reward obtained at $t$, and $\omega_{t}$ is the planned option provided by the LLM at $t$. The penalty factor $\lambda$ is used to balance the importance of avoiding unnecessary interactions. Note that if the decision made by the asking policy is “Not Ask” ($y_{t}==\textit{Not Ask}$), we set $\omega_{t}$ to be the plan executed at the previous timestep, namely let $\omega_{t}=\omega_{t-1}$. This ensures that if the agent decides not to ask for a new plan, it continues executing the same plan as before. During each iteration, data is collected on-policy using the model $\pi^{\text{ask}}_{\theta}$.
@@ 5 Experiments
We seek to address the following questions by experiments, : can our agent effectively reduce interaction costs while maintaining a high target task completion rate, compared with baseline methods? Can our agent proactively seek assistance from an LLM in exploratory environments? The experimental results indicate that the answer to both questions is yes. As a byproduct, we find that our approach is able to tolerate the imperfection of a crucial component, the translator in the mediator module, which is used to convert observed images into textual descriptions (refer details to the Appendix). We employ two versions of the Vicuna model (Vicuna-7b and Vicuna-13b) Touvron et al. (2023) as LLM planners.
@@ 5.1 Baselines
Hard-coded  The timing and conditions for requesting new instructions from LLMs are manually determined by human experts for each option Wang et al. (2023c). The agent will only request a new plan from the LLM planner when specific termination conditions for the option are met. These conditions involve a goal-finishing detector and a constraint on the maximum number of allowed timesteps. For example, let’s consider the option “go to the red door.” The termination condition for this option specifies that the agent should reach the target door location or exceed 100 timesteps spent on this option.
@@
Never  The agent never interacts with the LLM. Instead, the policy-over-options (i.e., the planner) is learned using RL techniques based on data collected during interactions with the environment Sutton et al. (1999); Precup (2000). This means that the agent learns to make decisions and generate plans without actively querying the LLM in real-time decision-making. By comparing this method with other approaches, we can assess the contribution of using an LLM as the planner. This comparison helps evaluate the effectiveness and advantages of incorporating a pre-trained LLM into the planning process.
@@ 5.2 MiniGrid Experiments
The MiniGrid environment consists of a collection of 2D grid-world environments with goal-oriented tasks Chevalier-Boisvert et al. (2023). In these environments, the agent must navigate within a 2D grid room and interact with specific objects to complete various tasks, such as “open the red door” or “put the green ball next to the yellow box”.
One important characteristic of this environment is that the agent’s view range is limited. This means that the agent needs to explore the environment and gather useful information to plan its actions effectively. The environment returns observations in the form of a full grid, but with unexplored areas occluded, similar to the concept of “fog of war” in games like StarCraft. Technically, the observation returned by the environment has a shape of $o\in\mathbb{R}^{W\times H\times 4}$, where $W$ and $H$ represent the width and height of the grid, respectively. For an unexplored grid at location $[w,h]$, the observation returns the vector $[-1,-1,-1,-1]$. For an explored grid, the corresponding 4D vector contains information about the object ID, color ID, state ID (e.g., closed or locked for a door), and the agent’s direction ID (indicating the agent’s orientation if it is present at this location, or 4 otherwise). This design allows us to focus on the agent’s reasoning ability and exclude potential influences from factors like memorization. Fig. 3 provides an example of the environment setup in the SimpleDoorKey scenario.
@@
Figure 4: An example of the prefix prompt and one interaction for the ColoredDoorKey environment. Prefix prompt consists of task instruction and few-shot examples. In Chain-of-Thought-style prompts, we add inference processes within the examples. Note that these few-shot examples are only provided for grounding a few but not all task-related knowledge to, and constraining the output formats, of the LLM. We do not need to exhaustively enumerate all knowledge and rules to construct prompts, as a qualified LLM can do logical reasoning based on a limited number of prompts, then provide proper plans (instructions) that are adaptable to new scenarios encountered in the environment.
In our experiments, we focus on the task of opening a locked door in five distinct environments: SimpleDoorKey, KeyInBox, RandomBoxKey, ColoredDoorKey, and MovingObstacle. All of these environments are procedurally generated, i.e., the grid layout (including room size, key and door locations) is randomly determined each time the environment is reset. To evaluate generalization, a held-out test set consisting of 100 randomly selected seeds is predefined for each environment. Refer to the Appendix for more details.
We use the Vicuna-7b model for the SimpleDoorKey, KeyInBox, RandomBoxKey, and MovingObstacle environments, while for the more complex ColoredDoorKey environment we use the Vicuna-13b model. As demonstrated in previous work Min et al. (2022), language models like LLMs require carefully designed prompts and few-shot demonstrations to generalize to different tasks. In our experiments, we provide task instructions and few-shot examples as in-context prompts for each environment. These prompts serve to guide the LLM to understand the task. For the challenging reasoning task in the ColoredDoorKey environment, we utilize Chain-of-Thought prompts proposed by Wei et al. (2022). These prompts help the LLM to deal with complex reasoning tasks specific to the ColoredDoorKey environment. The few-shot examples in the prompts are used to constraint the output formats. The LLM planner must utilize its ability to generalize and reason in order to comprehend the target task and adjust to situations that deviate from the few-shot examples, such as variations in objects’ colors. Fig. 4 provides an example of the prefix prompts and an interaction example in the ColoredDoorKey environment. It shows the effective performance of the LLM planner in producing an accurate plan in response to new observations.